Solve the linear system:
x-3y=-16,2x-7y=33

Answer by Maths68(1474)

Solve the linear system:
x-3y=-16,2x-7y=33
Solution
x-3y=-16 …………….1
2x-7y=33 ….………..2
Multiply equation 1 by -2 (to eliminate x)
-2x+6y=32 ……………3
Now add equation 3 and equation 2
-2x+6y=32
2x-7y=33
--------------------
-y=65 or y=-65
Put value of y in either equation 1 or 2 to get the value of x
x-3y=-16
x-3(-65)=-16
x+195=-16
x=-16-195
x=-211
(x,y)=(-211, -65)
Check your answer:-
Put values of x and y in either 1 or 2 equations
x-3y=-16
-211-3(-65)=-16
-211+195=-16
-16=-16

2x-7y=33
2(-211)-7(-65)=33
-422+455=33
33=33